75cc of water sample require 10.2cc of 0.01M EDTA for titration in presence of eriochrome black T indicator. Calculate the total hardness.
1
Expert's answer
2020-03-05T07:56:19-0500

The hardness of water is mainly due to the dissolved calcium salts, while magnesium is also presented in the drinking water. The interaction between EDTA and calcium cations:


"Ca^{2+}+EDTA^{4-} \rightarrow Ca(EDTA)^{2-};"


"C_{total}=\frac{C(EDTA)*V_{av}}{V_{w}}=\frac{0.01M*10.2mL}{75mL}=1.36 mmol/L;"


This could be converted:


"TotalHardness=1.36mmol\/L*40mg\/mmol=54.4mg\/L;"


or "54.4 \mu g/mL."